# RIGNO: A Graph-based framework for robust and accurate operator learning for PDEs on arbitrary domains

### v1

@@ 2 Methods
A problem with this approach is that the smallest prediction lead time $\tau_{min}$ in the training set is given by the minimum of $t_{n}-t_{n-1}$, for all $n$. Hence, at inference, predicting for lead times $\tau<\tau_{min}$ could lead to distribution shifts as smaller lead times have not been seen during training. Yet, a fundamental requirement for a time-dependent neural operator is to predict arbitrarily small times correctly in order to ensure time continuity. We tackle this problem by proposing to further fine-tune a RIGNO model, trained with the loss function (12), by using a modified loss function where $\widehat{u}_{m}^{t_{k}}({t_{n}})$ in (12) is obtained by starting from an intermediate input function: $\mathcal{G}_{\theta}(\widetilde{u}_{m}({t_{\ast}}),c_{m},t_{\ast},t_{n}-t_{\ast}$). Thus, our aim is to use solutions at intermediate time steps, which are not available in the training set, to make predictions at later time steps, available in the training data. However, we must find an approximation $\widetilde{u}_{m}({t_{\ast}})\approx u_{m}({t_{\ast}})$ as $u_{m}(t_{\ast})$ is not available during training. Our solution is to use the pretrained RIGNO itself by setting,
$\tilde{u}({t_{\star}})=\mathcal{G}_{\theta}\left(u(t_{k}),\;c,\;t_{k},\;t_{\star}-t_{k}\right).$
This self-consistent approach is reasonable as the modified loss function is only used during fine-tuning. Moreover, as shown in SM Figure C.3, the flow of the gradient can be stopped at the intermediate time $t_{\star}$, similar to ideas introduced in [4], which will be referred to as GradCut. The utility of this fractional pairing strategy will be tested subsequently.
@@
Table 1: Benchmarks on datasets with unstructured and uniform-grid space discretizations. Lowest (blue) and second lowest (orange) errors are highlighted with colors. The errors for time-dependent problems are calculated with three different autoregressive rollouts; and the lowest error is reported for each architecture. The details of autoregressive rollouts are available in SM Section C.2. All trainings are done with 1024 samples, except for the AF dataset where 2048 training samples are used. We observed that the relative errors were very sensitive to the normalization applied on the target values. More details about the metric used are available in SM Section E.
@@
(unstructured) RIGNO-18 RIGNO-12 GeoFNO FNO DSE GINO
Heat-L-Sines 0.04 0.05 0.15 0.53 0.19
Wave-C-Sines 5.35 6.25 13.1 5.52 5.82
NS-Gauss 2.29 3.80 41.1 38.4 13.1
NS-PwC 1.58 2.03 26.0 56.7 5.85
NS-SL 1.28 1.91 24.3 29.6 4.48
NS-SVS 0.56 0.73 9.75 26.0 1.19
CE-Gauss 6.90 7.44 42.1 30.8 25.1
CE-RP 3.98 4.92 18.4 27.7 12.3
ACE 0.01 0.01 1.09 1.29 3.33
Wave-Layer 6.77 9.01 11.1 28.3 19.2
AF 1.00 1.09 4.48 1.99 2.00
Elasticity 4.31 4.63 5.53 4.81 4.38
(uniform grid) RIGNO-18 RIGNO-12 CNO scOT FNO
@@
3 Results
To test the performance of RIGNO for PDEs on arbitrary domains, we choose two types of datasets. First, we consider 4 datasets on structured and unstructured meshes, namely Heat-L-Sines (heat equation on a L-shaped domain), Wave-C-Sines (wave equation, propagation of waves in a homogeneous medium of circular shape), AF (compressible transonic flow past airfoils taken from [21]), and Elasticity (hyper-elastic deformations of varying domain shapes, also taken from [21]). See SM Section D for details of datasets and their abbreviations used here. Next, we test how well RIGNO learns operators defined on arbitrary point clouds. To set this up, we choose 8 datasets from the PDEGym database [12], which deal with four different PDEs; the incompressible Navier-Stokes equations, the compressible Euler equations, the Allen-Cahn equation, and the wave equation on a spatially varying medium. The datasets are all originally defined on uniform Cartesian grids. For the unstructured experiments, we randomly subsample 9216 points from the available coordinates on a 128$\times$128 uniform grid, see SM Figure D.1.
In Table 1, we present the test errors on these 12 unstructured datasets [31] with two sizes of RIGNO, namely RIGNO-12 and RIGNO-18, with 12 and 18 layers in their processor modules, respectively. To contextualize the results, we consider three baselines, namely Geo-FNO [21], FNO DSE [23] and GINO [22]. All of the baselines are capable of handling inputs and outputs on arbitrary point clouds, yet use different routes to achieve this. More details about the baselines, including the selection of hyperparameters, are outlined in SM Section F. We observe from Table 1 that RIGNO-18 outperforms all the baselines in all 12 datasets. Likewise, RIGNO-12 is the second best-performing model for 10 of the 12 datasets. For Wave-C-Sines and Elasticity, FNO DSE and GINO respectively claim second best, yet RIGNO-12 remains comparable in performance to RIGNO-18. Moreover, in some cases, the RIGNO variants are almost an order of magnitude more accurate than the baselines, showing excellent performance on these benchmarks. We would also like to emphasize that RIGNO is the only framework that displays acceptably low errors for every single dataset that we consider here, demonstrating its robustness.
Often, models are designed to have flexible inputs/outputs by potentially reducing their performance on Cartesian domains. To further illustrate the expressive power of RIGNO, we consider the same 8 datasets from the PDEGym database as before, but on their original Cartesian grids. In this case, we can compare RIGNO with state-of-the-art neural operators such as FNO, CNO, and scOT which are tailored for Cartesian grids. The resulting test errors, presented in Table 1, also show that RIGNO is very accurate on Cartesian grids and RIGNO-18 is the best performing model on every single dataset. Again, RIGNO-12 is either the second best-performing model or comparable to it. In fact, for most datasets, RIGNO-18 is several times more accurate than neural operators such as CNO and scOT, as well as the widely used FNO, showcasing the expressive power of its GNN-based backbone.
@@ 3 Results
(b) Model size.
Figure 2: Autoregressive test errors at $t=t_{14}$. Fractional pairing is off in all trainings. All results correspond to the AR-2 rollout scheme (defined in SM Section C.2). All axes are in logarithmic scales. The plot on the right corresponds to the NS-PwC dataset.
RIGNO scales with data and model size.
An essential requirement for modern ML models is their ability to scale with data, i.e., the test error should reduce at a consistent rate as the number of training samples increases [13, 17, 12]. To test this aspect, we plot the test error with respect to the number of training samples in Figure 2(a), observing that RIGNO-12 shows a consistent scaling for the datasets. Another critical aspect of ML models is their ability to scale with model size [13, 12]. Compared to data scaling, scaling with model size is more subtle as a number of hyperparameters influence model size and changing each of them, while keeping the others fixed, might lead to different outcomes. In SM Section G.2, we present a detailed discussion and results on how RIGNO scales with respect to each of the hyperparameters that influences model size. Here, in Figure 2(b), we only consider two (correlated) indications of model size, i.e., total number of model parameters and inference time. We see from Figure 2(b) that RIGNO-12 consistently scales with respect to both the total number of parameters and inference run time.
@@
(c) Time-stepping strategies (NS-SL).
Figure 3: Ablation of the proposed training strategies. a) The effect of edge masking on resolution invariance. EMP stands for edge masking probability. The y-axis shows the relative test error (at $t=t_{14}$) when the spatial resolution of the inputs and outputs varies from the training resolution ($92^{2}$). b) Autoregressive test errors ($\tau=\Delta t$) with and without fractional pairing fine-tuning. All values after the vertical line are considered as extrapolation in time. c) Autoregressive test errors ($\tau=2\Delta t$) with different time-stepping methods. All models are trained without the fractional pairing strategy and with $\tau_{\rm max}=8\Delta t$. All values after the vertical lines are considered as extrapolation in time.
@@
RIGNO is designed to learn the solution operators of time-dependent PDEs and contains many elements in its design to enforce time continuity and enable accurate temporal approximation. To test these aspects, we focus on the NS-PwC and the NS-SL datasets (see SM Section D). The underlying trajectories are available from the PDEgym collection [12] on a time interval of $[0,1]$, sampled at equidistant time steps $t_{n}=n\Delta t$, with $\Delta t=0.05$ and $0\leq n\leq 20$. We trained RIGNO with data sampled at timesteps $\{t_{0},t_{2},t_{4},t_{6},t_{8},t_{10},t_{12},t_{14}\}$. Thus, we set $\tau=2\Delta t$ in our temporal marching procedure. During inference, we unroll RIGNO autoregressively with a time step size of $\Delta t$ and compare the test errors at all possible time indices $1\leq n\leq 20$. The resulting test errors are shown in Figure 3 (b,c). We have several observations to make about the results. i) The test errors grow in time as expected from an autoregressive rollout strategy. However, the growth of errors is not exponential but rather linear. ii) RIGNO is able to extrapolate well outside the temporal training domain. We see this from comparing the errors at time indices $n\geq 15$. These times have not been seen during training. Yet, RIGNO is able to generalize very well even to these unseen time points. iii) The proposed fractional pairing fine-tuning approach significantly enhances the accuracy with small unseen time steps. The errors are reduced by a significant factor, and in particular, fractional pairing helps generalize much better in the time extrapolation regime, $n\geq 15$, with significantly lower errors and lower growth rate of errors as compared to when it is switched off.
We also consider time marching strategies arising from the choice of parameters $\alpha,\beta$ in (11). We present the results with all the three choices of time stepping, i.e., output, residual and derivative stepping for the NS-SL dataset in Figure 3 (c). In addition to the slow error growth and good extrapolation beyond the times seen during training, we see that the derivative time-stepping strategy with $\alpha=1,\beta=\tau$ is clearly superior in this example to the other two, leading us to prefer it during our evaluations.
Resolution Invariance and RIGNO.
As mentioned before, a key requirement for a robust neural operator is to enforce an appropriate form of resolution invariance [2]. To test the abilities of RIGNO in this regard, we consider the NS-PwC dataset on random point clouds (see SM Section D). The training resolution for RIGNO is $92^{2}$. At inference, we test it on this and several other resolutions, presenting the test errors in Figure 3(a). We see from this figure that the resolution invariance errors with the base version of RIGNO are fairly small and indicate that the proposed local aggregation scheme enables RIGNO to generalize well across spatial resolution. Additionally, we observe that edge masking allows us to enforce resolution invariance to a greater extent. Moreover, the computational cost is roughly cut to half with edge masking probability $0.5$. As shown in Figure 2(b) and SM Figure G.1, using a larger model can compensate for the accuracy drop caused by edge masking.
@@
Not only does edge masking enable RIGNO to be (approximately) resolution invariant, it provides a simple method to quantify model uncertainties. As edge masking is random, invoking it during inference enables the model to provide multiple outputs for the same input. The mean of the outputs can be used as the model prediction as shown in SM Figure G.3. The same figure shows how the standard deviation of this ensemble of outputs is highly correlated with the test error. Thus, it can be used to provide an adaptive time marching strategy to reduce errors at inference, see SM Section G.4 for details. In SM Section G.5, we show that edge masking makes RIGNO more robust to noisy inputs as well.
4 Discussion.
@@
In addition to this basic construction (see Figure 1), which was inspired by [16], RIGNO introduces many novel elements to satisfy two important requirements. First, as desired in [15, 2], (approximate) resolution invariance is essential for structure-preserving operator learning. RIGNO uses overlaps between the sub-regions in the regional mesh connecting points in the physical mesh and more importantly, edge masking, to enable approximate resolution invariance, a feature which is hard to achieve for neural operators on arbitrary domains. Second, RIGNO is tailored to process time-dependent PDEs. In this context, temporal continuity and control on error growth over time are key properties that the neural operator needs to uphold. Here, we have proposed many elements in RIGNO to achieve these requirements, namely lead time conditioning, all2all training, and a fractional pairing fine-tuning strategy. This allows us to maximize training steps for a given number of trajectories, while allowing for general rollout strategies at inference.
@@
We have tested RIGNO on a suite of challenging problems including PDEs with data on structured/unstructured meshes, random point clouds, and Cartesian grids. RIGNO is shown to be significantly more accurate than baselines designed for such arbitrary domains. We also observe that RIGNO is more accurate than state-of-the-art neural operators such as CNO, FNO and scOT, even for PDEs on Cartesian grids. The gains in accuracy are not marginal, and can be of an order of magnitude on many problems, demonstrating the expressive power of the proposed architecture. We show that RIGNO scales with both model and data size. Simultaneously, RIGNO has excellent resolution invariance and temporal continuity properties. It also generalizes very well to interpolated and extrapolated time levels, validating novel elements such as fractional pairing that were introduced into the model to improve temporal processing.
@@
Importantly, we must also highlight some limitations of RIGNO and indicate avenues for further extension. Here, we have focused on PDEs in two space dimensions. In principle, RIGNO can readily be extended to PDEs in three space dimensions, although the associated computational cost and paucity of publicly available 3D benchmarks currently limits our exploration of this crucial direction. Similarly, our focus here was on presenting the model and extensive empirical investigation. Important theoretical questions such as those of universal approximation of PDE solution operators with RIGNO will be considered in future work. Similarly, we have empirically demonstrated that RIGNO scales with the number of training samples; however, proving a rigorous generalization error bound, such as that of [17], merits further analysis. Moreover, the scaling curves of Figure 2 suggest that a large number of training samples are necessary for neural operators to achieve small error. This issue is common to all standalone neural operators [12] and can be alleviated by building foundation models such as those proposed in [12, 11, 41, 29]. Using RIGNO as the backbone for a foundation model is another avenue for further exploration. Finally, downstream applications of RIGNO to tasks such as uncertainty quantification (UQ), inverse problems and PDE-constrained optimization will be considered in future work.

### v2

@@ 2 Methods
A problem with this approach is that the smallest prediction lead time $\tau_{min}$ in the training set is given by the minimum of $t_{n}-t_{n-1}$ for all $n$. Hence, at inference, predicting for lead times $\tau<\tau_{min}$ could lead to distribution shifts as smaller lead times have not been seen during training. Yet, a fundamental requirement for a time-dependent neural operator is to predict arbitrarily small times correctly in order to ensure time continuity. We tackle this problem by proposing to further fine-tune a RIGNO model, trained with the loss function (4), by using a modified loss function where $\widehat{u}_{m}^{t_{k}}({t_{n}})$ in (4) is obtained by starting from an intermediate input function: $\mathcal{G}_{\theta}(\widetilde{u}_{m}({t_{\star}}),c_{m},t_{\star},t_{n}-t_{\star}$). Thus, our aim is to use solutions at intermediate time steps, which are not available in the training set, to make predictions at later time steps, available in the training data. However, we must find an approximation $\widetilde{u}_{m}({t_{\star}})\approx u_{m}({t_{\star}})$ as $u_{m}(t_{\star})$ is not available during training. Our solution is to use the pretrained RIGNO itself by setting $\tilde{u}({t_{\star}})=\mathcal{G}_{\theta}\left(u(t_{k}),\;c,\;t_{k},\;t_{\star}-t_{k}\right)$. This self-consistent approach is reasonable as the modified loss function is only used during fine-tuning. Recycling model outputs as inputs has been employed in previous implementations for the fixed-step setting [4, 20] to improve rollout stability. Here, we extend this concept to the time-continuous setting, aiming to generalize to arbitrarily small time steps. As shown in SM Figure C.3, the flow of the gradient can be stopped at the intermediate time $t_{\star}$ for training efficiency. The utility of this fractional pairing strategy will be tested subsequently.
@@
3 Results
To test the performance of RIGNO for PDEs on arbitrary domains, we choose two types of datasets. First, we consider 4 datasets on point clouds obtained from structured and unstructured meshes, namely Heat-L-Sines (heat equation on a L-shaped domain), Wave-C-Sines (wave equation, propagation of waves in a homogeneous medium of circular shape), AF (compressible transonic flow past airfoils taken from [19]), and Elasticity (hyper-elastic deformations of varying domain shapes, also taken from [19]). Next, we test how well RIGNO learns operators defined on random point clouds. To set this up, we choose 9 datasets from the PDEGym database [13], which deal with four different PDEs; the incompressible Navier-Stokes equations, the compressible Euler equations, the Allen-Cahn equation, the wave equation on a spatially varying medium, and the Poisson equation. The datasets are all originally defined on uniform Cartesian grids. To obtain a random point cloud, we randomly subsample 9216 points from the available coordinates on a $\times128$ uniform grid, see SM Figure D.1. Our dataset collection spans 8 PDEs and comprises 3 time-independent and 10 time-dependent problems. See SM Section D for more details about the datasets and their abbreviations used here.
In Table 1, we present the test errors on these 13 unstructured datasets [32] with two sizes of RIGNO, namely RIGNO-12 and RIGNO-18, with 12 and 18 layers in their processor modules, respectively. To contextualize the results, we consider five baselines, namely MeshGraphNet [33], Geo-FNO [19], FNO DSE [24], GINO [23], and UPT [33]. All of these baselines are capable of handling inputs and outputs on arbitrary point clouds, yet use different routes to achieve this. More details about the baselines, including the selection of hyperparameters, are outlined in SM Section G. We observe from Table 1 that RIGNO-18 outperforms all the baselines in all 13 datasets. Likewise, RIGNO-12 is the second best-performing model for 10 of the 13 datasets. For Wave-C-Sines, Poisson-Gauss, and Elasticity, other architectures claim second best; yet, RIGNO-12 remains comparable in performance to RIGNO-18. Moreover, in some cases, the RIGNO variants are almost an order of magnitude more accurate than the baselines, showing excellent performance on these benchmarks.
Table 1: Benchmarks on datasets with unstructured and Cartesian-grid space discretizations. Lowest and second lowest errors are highlighted. The errors for time-dependent problems are the lowest among three autoregressive strategies (see SM Section C.2). All training datasets have 1024 trajectories/samples except for the AF dataset (2048). MGN corresponds to the MeshGraphNet baseline.
@@ 3 Results
(point cloud) RIGNO-18 RIGNO-12 MGN Geo-FNO FNO DSE GINO UPT
Heat-L-Sines 0.04 0.05 0.51 0.15 0.53 0.19 20.2
Wave-C-Sines 5.35 6.25 17.3 13.1 5.52 5.82 12.7
NS-Gauss 2.29 3.80 76.4 41.1 38.4 13.1 92.5
NS-PwC 1.58 2.03 33.1 26.0 56.7 5.85 100
NS-SL 1.28 1.91 18.8 24.3 29.6 4.48 51.5
NS-SVS 0.56 0.73 7.53 9.75 26.0 1.19 4.2
CE-Gauss 6.90 7.44 76.4 42.1 30.8 25.1 64.2
CE-RP 3.98 4.92 13.9 18.4 27.7 12.3 26.8
ACE 0.01 0.01 0.09 1.09 1.29 3.33 100
Wave-Layer 6.77 9.01 84.8 11.1 28.3 19.2 19.6
Poisson-Gauss 2.26 2.52 30.9 8.16 2.27 7.57 48.4
AF 1.00 1.09 10.1 4.48 1.99 2.00 45.7
Elasticity 4.31 4.63 11.9 5.53 4.81 4.38 12.6
Table 2: Benchmarks on datasets with Cartesian-grid space discretizations. Lowest and second lowest errors are highlighted. The errors for time-dependent problems are the lowest among three autoregressive strategies (see SM Section C.2). All training datasets have 1024 trajectories/samples.
Dataset Median relative $L^{1}$ error [%]
(Cartesian grid) RIGNO-18 RIGNO-12 CNO scOT FNO
@@
Poisson-Gauss 1.80 2.44 2.04 0.68 11.46
Often, models are designed to have flexible inputs/outputs by potentially reducing their performance on Cartesian domains. To further illustrate the expressive power of RIGNO, we consider 9 datasets from the PDEGym database on their original Cartesian grids. In this case, we can compare RIGNO with state-of-the-art neural operators such as FNO, CNO, and scOT, which are tailored for Cartesian grids. The resulting test errors in Table 2 show that RIGNO is very accurate on Cartesian grids too and RIGNO-18 is the best performing model on 8 out of 9 datasets, often with several times lower errors than the baselines. We would also like to emphasize that RIGNO is the only framework that displays acceptably low errors for every single dataset that we consider here, demonstrating its unprecedented robustness.
RIGNO scales with data and model size.
An essential requirement for modern ML models is their ability to scale with data, i.e., the test error should reduce at a consistent rate as the number of training samples increases [14, 18, 13]. To test this aspect, we plot the test error with respect to the number of training samples in Figure 2(a), observing that RIGNO-12 shows a consistent scaling for the datasets. Another critical aspect of ML models is their ability to scale with model size [14, 13]. Compared to data scaling, scaling with model size is more subtle as a number of hyperparameters influence model size and changing each of them, while keeping the others fixed, might lead to different outcomes. In SM Section F.3, we present a detailed discussion and results on how RIGNO scales with respect to each of the hyperparameters that influences model size. Here, in Figure 2(b), we only consider two (correlated) indications of model size, i.e., total number of model parameters and inference time. We see from Figure 2(b) that RIGNO-12 consistently scales with respect to both the total number of parameters and inference run time.
@@
(b) Model size (NS-PwC).
Figure 2: Scaling behavior of RIGNO with respect to dataset and model size. All values correspond to autoregressive (AR-2, see SM Section C.2) test errors at $t=t_{14}$.
Resolution invariance and RIGNO.
As mentioned before, a key requirement for a robust neural operator is to enforce an appropriate form of resolution invariance [2]. To test the abilities of RIGNO in this regard, we consider the NS-PwC dataset on random point clouds. The training resolution for RIGNO is $42^{2}$. At inference, we test it on this and several other resolutions, presenting the test errors in Figure 3(a). We see from this figure that the super-resolution test errors with the base version of RIGNO are very small and indicate that the proposed local aggregation scheme enables RIGNO to generalize well across higher spatial resolutions. The results of the same test on MeshGraphNet (see SM Section F.4) provide further evidence that local aggregations are essential for obtaining super-resolution invariance. Furthermore, the proposed edge masking technique allows us to enforce resolution invariance to a greater extent, particularly in the sub-resolution regime. Moreover, we observe that edge masking substantially improves the optimization process by allowing the model to learn weaker spatial correlations, while simultaneously cutting the computational cost by a factor of 2 when $50\%$ of the edges are masked.
@@
(c) Time marching (NS-SL).
Figure 3: Ablation of the proposed training strategies. a) The effect of edge masking on resolution invariance. EM stands for edge masking. The y-axis shows the relative test error at $t=t_{14}$ when the spatial resolution of the inputs and outputs varies from the training resolution $64^{2}$. b) Autoregressive test errors with and without fractional pairing fine-tuning. The initial solution (model input) is updated on vertical lines according to the AR-4 scheme (see SM Section C.2). These models are exceptionally trained on snapshots with time resolution $4\Delta t$ and up to time $t_{16}$. c) Autoregressive (AR-2) test errors with different time marching strategies. Time snapshots after the vertical line has not been seen during training and are considered as extrapolation in time.
@@
RIGNO can learn the solution operators of time-dependent PDEs and contains several elements in its design to enforce time continuity and enable accurate temporal approximation. To test these aspects, we focus on the NS-PwC and the NS-SL datasets. The underlying trajectories are available from the PDEgym collection [13] on a time interval of $[0,1]$, sampled at equidistant times $t_{n}=n\Delta t$, with $\Delta t=0.05$ and $0\leq n\leq 20$. We trained RIGNO with data sampled at times $\{t_{0},t_{4},t_{8},t_{12},t_{16}\}$. The smallest time difference present in the training dataset is thus $\tau=4\Delta t$. During inference, we unroll RIGNO using time steps $\tau\in\{\Delta t,2\Delta t,3\Delta t,4\Delta t\}$ according to the AR-4 scheme described in SM Section C.2. We compare the test errors at all possible time indices $1\leq n\leq 20$. Hence, we test the generalization capability of RIGNO to a higher time resolution. The resulting test errors are shown in Figure 3(b). With regular training, the errors with unseen time steps $\tau<4\Delta t$ are significantly higher than with $\tau=4\Delta t$, leading to the observed jagged pattern in the plot of the error over time. The proposed fractional pairing fine-tuning approach effectively enhances the accuracy with small unseen time steps, leading to a monotone error accumulation with $t\geq t_{4}$, where the output time is in-distribution.
We also consider time marching strategies arising from the choice of parameters $\alpha,\beta$ in (3). Here, we train and infer the model using the same time resolution ($2\Delta t$). We present the results with all the three choices of time marching strategies, namely output, residual and derivative marching for the NS-SL dataset in Figure 3(c). We observe a slow error growth and very good extrapolation beyond the times seen during training ($t>t_{14}$). Moreover, we see that the derivative time marching strategy with $\alpha=1,\beta=\tau$ is clearly superior in this example to the other two, leading us to prefer it during our evaluations.
@@
Not only does edge masking enable RIGNO to be (approximately) resolution invariant, but it also provides a simple method to quantify model uncertainties. As edge masking is random, invoking it during inference enables the model to provide multiple outputs for the same input. The mean of the outputs can be used as the model prediction as shown in SM Figure F.4. The same figure shows how the standard deviation of this ensemble of outputs is highly correlated with the test error. Thus, it can be used to provide an adaptive time marching strategy to reduce errors at inference, see SM Section F.6 for details. In SM Section F.7, we show that edge masking makes RIGNO more robust to noisy inputs as well.
4 Discussion
@@
In addition to this basic construction (see Figure 1), RIGNO introduces several novel elements to satisfy two important requirements. First, as desired in [16, 2], (approximate) resolution invariance is essential for structure-preserving operator learning. RIGNO uses overlaps between the sub-regions in the regional mesh connecting points in the physical mesh and more importantly, edge masking, to enable approximate resolution invariance, a feature which is hard to achieve for neural operators on arbitrary domains. Second, RIGNO is tailored to process time-dependent PDEs. In this context, temporal continuity and control on error growth over time are key properties that the neural operator needs to uphold. Here, we have incorporated many elements in RIGNO to achieve these requirements, namely lead time conditioning, all2all training, and a novel fractional pairing fine-tuning strategy. This allows us to maximize training steps for a given number of trajectories, while allowing for general rollout strategies with arbitrarily small time steps at inference.
@@
We have tested RIGNO on a suite of challenging problems including PDEs with data on structured/unstructured meshes, random point clouds, and Cartesian grids. RIGNO is shown to be significantly more accurate than baselines designed for such arbitrary domains. We also observe that RIGNO is more accurate than state-of-the-art neural operators such as CNO, FNO and scOT, even for PDEs on Cartesian grids. The gains in accuracy are not marginal, and can be of an order of magnitude on many problems, demonstrating the expressive power of the proposed architecture. We attribute RIGNO’s state-of-the-art accuracy and robustness primarily to the high expressive power of local message passing, which enables fine-grained manipulation and propagation of information. This advantage is especially prominent in time-dependent and advection-dominated problems. For elliptic problems, where global interactions are more crucial, the gains are less pronounced. To mitigate the limitations of purely local interactions, RIGNO incorporates multi-scale edges in the processor, introducing long-range interactions that significantly improve accuracy without incurring notable additional computational cost. Furthermore, we show that RIGNO scales with both model and data size, while having excellent resolution invariance and temporal continuity properties. It generalizes very well to interpolated and extrapolated time levels, and maintains its accuracy with very small time steps.
Related work.
The popular MeshGraphNet [33] served as a foundation for the formulation of RIGNO and thus admits basic architectural similarities. This approach constructs a graph on the input coordinates based on a mesh and performs message passing steps directly on the physical coordinates. Yet, as confirmed by our empirical results (see SM Section F.4), this approach leads to a strong dependence on the space resolution of the training data and cannot generalize to a different resolution. Furthermore, the computational cost quickly becomes prohibitive as the number of mesh (input) points increases, hindering MeshGraphNet from scaling to larger problems. This issue is addressed in [10] by downsampling the initial mesh to a coarser mesh and performing more message passing steps on the coarse mesh. However, this approach still relies on the original mesh for resolving small scales and consequently inherits the dependence on the training resolution from MeshGraphNet. As evidenced by our experimental results, the multi-scale regional mesh construction of RIGNO mitigates the reliance of such graph-based approaches on the distribution of the input points and introduces the ability to scale to larger problem sizes (see Table F.2). Radius-based aggregation for projection onto a coarse grid has been explored in prior work and shown to be effective for non-graph-based architectures. GINO [23] performs local aggregation of coordinates within a fixed radius to map between an unstructured point cloud and a uniform mesh, enabling the use of FNO as the core processor. UPT [1] employs transformers as the core processor, and thus does not require the mapped coordinates to lie on a uniform grid. In this case, local aggregation is optionally applied in the encoder to compress the information from dense point clouds to a coarser representation. Both these methods rely on a global fixed radius which is a hyperparameter and must be calibrated for each dataset. In practice, the given point clouds are often more refined in regions with sharper gradients, making a fixed-radius approach destined to produce an uneven number of edges per receiver node. To address this, additional mechanisms such as limiting the number of incoming edges per node are necessary. By contrast, the adaptive strategy used in the current work determines aggregation radii based on local resolution, thereby mitigating both these issues while automatically ensuring full domain coverage without the need for manual calibration.
Limitations.
@@
This work focuses on PDEs in two spatial dimensions. However, many of the considered datasets are comparable in complexity, grid resolution, and number of input-output pairs to typical 3D problems. In principle, RIGNO is readily extendable to PDEs in three dimensions, though the current paucity of publicly available 3D benchmarks limits our ability to explore this important direction. Furthermore, our focus here was on presenting the model and extensive empirical investigation. Important theoretical questions such as those of universal approximation of PDE solution operators with RIGNO will be considered in future work. Similarly, we have empirically demonstrated that RIGNO scales with the number of training samples; however, proving a rigorous generalization error bound, such as that of [18], merits further analysis. Moreover, the scaling curves of Figure 2 suggest that a large number of training samples are necessary for neural operators to achieve small error. This issue is common to all standalone neural operators [13] and can be alleviated by building foundation models such as those proposed in [13, 11, 44, 30]. Our results presented in SM Section F.9 show that RIGNO can largely benefit from transfer learning. Using RIGNO as the backbone for a foundation model is another avenue for further exploration. Finally, downstream applications of RIGNO to tasks such as uncertainty quantification (UQ), inverse problems and PDE-constrained optimization will be considered in future work.